OpenStudy (anonymous):

the sum of the length and girth of a rectangular box must be smaller than 96 in if the girth is 1/2 of the length,what is the length?

OpenStudy (anonymous):

64 inches

OpenStudy (anonymous):

x+0.5x = 96 then solve :)

OpenStudy (turingtest):

smaller-than, so solve\[\ell+g<96\to\ell+\frac{\ell}{2}<96\]

OpenStudy (anonymous):

lost on both

OpenStudy (anonymous):

ok i add

OpenStudy (anonymous):

i got 64.5

OpenStudy (turingtest):

let the length be l let the girth be g we are told that the sum of the length and girth must be "less-then" 96, so\[\ell+g<96\]we are also told that the girth is 1/2 the length, so\[\frac{\ell}{2}=g\]sub this into the expression above and we get\[\ell+\frac{\ell}{2}<96\]

OpenStudy (turingtest):

\[\ell+\frac{\ell}{2}=\frac{3\ell}{2}<96\]which leads to\[\ell<\frac{2}{3}\times96\]so what do you get for length?

OpenStudy (anonymous):

64.5?

OpenStudy (anonymous):

am i right plz

OpenStudy (turingtest):

I just can;t imagine where you get the .5 from...\[\frac{2}{3}\times96=64\](exact) so what did you do?

OpenStudy (anonymous):

the calculator again:(

OpenStudy (anonymous):

ok i multiplied 2/3 by 96 an got 64

OpenStudy (anonymous):

thx :)

OpenStudy (turingtest):

do the pieces separately \[96\times2=192\]\[\frac{192}{3}=64\]but remember that this is talking about our expression\[\frac{3\ell}{2}<96\]so the solution is\[\ell<64\]it is important to keep the less-than sign, it's not an equality!

OpenStudy (anonymous):

can u check 1 more problem for me i answered it but not sure if im right its the 1 about the area rug?

OpenStudy (turingtest):

sure thing :)